OpenStudy (anonymous):

simplify:( 3x+1/x+2)-(-5/x+7)

jhonyy9 (jhonyy9):

3x+1 -5 (3x+1)(x+7) +5(x+2) 3x^2 +22x +7 +5x +10 ----- - ------ = ------------------- = ---------------------- = x+2 x+7 (x+2)(x+7) (x+2)(x+7) 3x^2 +27x +17 = --------------- (x+2)(x+7)

OpenStudy (anonymous):

Need help walking through this one?

OpenStudy (anonymous):

yep im stuck yet again